OpenStudy (anonymous):

how do you find the zero(s) of the function f(x)=x^2-x+(5/4)

OpenStudy (bahrom7893):

4x^2 - 4x + 5 = 0

OpenStudy (anonymous):

why would you do that? common denominator?

OpenStudy (bahrom7893):

no because it looks a little prettier.. just multiply out by 4

OpenStudy (bahrom7893):

now use the discriminant formula

OpenStudy (anonymous):

discriminant formula..? feeling braindead tonight, ahhh

OpenStudy (anonymous):

-b/2a? no..

OpenStudy (anonymous):

that's slope/axisofsymmetry

OpenStudy (bahrom7893):

D = b^2-4ac x1 = (-b-sqrt(D))/(2a) x2 = (-b+sqrt(D))/(2a)

OpenStudy (anonymous):

the thing is though.. I have to sketch a graph.. how would I sketch those x's?

OpenStudy (anonymous):

parabola that opens up

OpenStudy (anonymous):

well yeah, but the x values, I have to designate a coordinate within as close proximity as possible..